There are two identical vessels filled with equal amount of ice.The vessels are of different metals,If the ice melts in the two vessels in 20 and 35 minutes respectively, the ratio of the coefficients of thermal conductivity of the two metals is

Grade:12th pass

1 Answers

Arun
6 years ago
Dear student
 
As we know
Q/t = K A \DeltaT/ L
For two identical vessels
k1 * t1 = k2 *t2
K1/k2 = t2/t1 = 35/20 = 7/4
Regards
Arun (askIITians forum expert)